FreshSites Limited Status
All Systems Operational
Front-End Services Operational
90 days ago
100.0 % uptime
Today
cloud.freshsites.co.uk ? Operational
90 days ago
100.0 % uptime
Today
freshsites.co.uk ? Operational
90 days ago
100.0 % uptime
Today
Managed Hosting Operational
LON1 ? Operational
FRA1 ? Operational
AMS1 ? Operational
NYC1 ? Operational
SFO1 ? Operational
SGP1 ? Operational
Standard Hosting Operational
fs01.freshsites.co.uk ? Operational
fs02.freshsites.co.uk ? Operational
fs03.freshsites.co.uk ? Operational
fs04.freshsites.co.uk ? Operational
fs05.freshsites.co.uk ? Operational
fs06.freshsites.co.uk ? Operational
fs12.freshsites.co.uk ? Operational
fs13.freshsites.co.uk ? Operational
fs14.freshsites.co.uk ? Operational
fs15.freshsites.co.uk ? Operational
fs16.freshsites.co.uk ? Operational
fs17.freshsites.co.uk ? Operational
fs18.freshsites.co.uk ? Operational
fs19.freshsites.co.uk ? Operational
fs20.freshsites.co.uk ? Operational
fs21.freshsites.co.uk ? Operational
Email Operational
Email ? Operational
Backup Operational
Backup Ireland Operational
Backup London Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
had a major outage
had a partial outage
Past Incidents
May 23, 2019

No incidents reported today.

May 22, 2019

No incidents reported.

May 21, 2019

No incidents reported.

May 20, 2019

No incidents reported.

May 19, 2019

No incidents reported.

May 18, 2019

No incidents reported.

May 17, 2019

No incidents reported.

May 16, 2019

No incidents reported.

May 15, 2019

No incidents reported.

May 14, 2019
Postmortem - Read details
May 15, 10:15 BST
Resolved - This incident has been resolved.
May 14, 17:41 BST
Monitoring - This mornings issue has been identified on the server. The Swap on the disk had maxed out, which in turn caused the migration process (this manages job swapping between the server CPUs) to cause the server to hang.

This has been found to be a bug in the server Kernel which has now been upgraded, with a further server reboot being required for out of hours on 15/05/2019.
May 14, 11:22 BST
Investigating - We are currently investigating issues with server FS21. Customers on this server may currently experience site and email connectivity issues.
May 14, 10:13 BST
May 13, 2019

No incidents reported.

May 12, 2019

No incidents reported.

May 11, 2019

No incidents reported.

May 10, 2019

No incidents reported.

May 9, 2019

No incidents reported.